Calculate the boiling point of a 0.2 m aqueous sugar, C12H22O11 in ethanol.


1
Expert's answer
2021-05-19T02:29:01-0400

Kb (ethanol) = 1.23

Tb (ethanol) = 78.37 oC.

Tb (solution) = 78.37 + 1.23"\times"0.2 = 78.62 oC.

Answer: 78.62 oC.